OpenStudy (perii224):

Slopes? Find the slope of the line passing through points A(1, -2), F(4, -8).

OpenStudy (anonymous):

You need the formula for slope...

OpenStudy (anonymous):

(x1, y1) and (x2, y2)

OpenStudy (anonymous):

You can plug the 2 points into those ^^

OpenStudy (anonymous):

It doesn't really matter which ones go where.

OpenStudy (perii224):

These Were The Answers.. 2 -2 ½ -1/2

OpenStudy (igreen):

\(m = \dfrac{y_2-y_1}{x_2-x_1}\) Plug them in: \(m = \dfrac{-8+2}{4-1}\) Can you solve that?

OpenStudy (igreen):

@perii224 Add & Subtract, then divide.

OpenStudy (anonymous):

y2 - y1 -8 + -2 -6 _______ = _________ ___ = ??? x2 - x1 4 - 1 3

OpenStudy (igreen):

It's not -8 + -2.

OpenStudy (anonymous):

-6 / 3 = ? @perii224

OpenStudy (perii224):

So I'm Guessing It's -2?

OpenStudy (igreen):

Yes, you got it.

OpenStudy (perii224):

Thank You! :D

OpenStudy (igreen):

\(\dfrac{-8 + 2}{4 - 1} \rightarrow \dfrac{-6}{3} \rightarrow -2\)

OpenStudy (anonymous):

Typo....

OpenStudy (anonymous):

You've never seen one before ?